OpenStudy (mony01):

Find the area of the region bounded by the parabola y = 5x^2, the tangent line to this parabola at (4, 80), and the x-axis.

terenzreignz (terenzreignz):

You'd do well to first find the equation of the aforementioned tangent line ^_^

OpenStudy (mony01):

is the equation y=40x-80

OpenStudy (anonymous):

Yes, it seems right to me.

OpenStudy (mony01):

then what do you do to get the area? I dont know how to set it up

terenzreignz (terenzreignz):

Try to sketch it, maybe it'll give you an idea...

terenzreignz (terenzreignz):

Crud... the draw option doesn't work... >.>

terenzreignz (terenzreignz):

Here's a rough sketch:

terenzreignz (terenzreignz):

We need to find the area of that yellow bit... catch me so far?

OpenStudy (mony01):

yea

terenzreignz (terenzreignz):

Okay, so this is an integration problem, one that will have to be divided into two parts because of the nature of the area that you need to, well, find the area of. First find the area of this shaded part, because it's the more straightforward integral:

OpenStudy (anonymous):

I believe, it can be done in one integral, if we take slices parallel to the x-axis and integrate with respect to y. There is nothing wrong with doing it the way @terenzreignz suggested. May be both take the same amount of time.

terenzreignz (terenzreignz):

That is... true :3 But I'm allergic to radicals XD

OpenStudy (anonymous):

You only have to deal with \( \sqrt y \)

OpenStudy (mony01):

how would i set the integral up?

terenzreignz (terenzreignz):

Well, going on from the method I put up, the area of the black shaded part is simply the integral of the parabola (quadratic) The only tricky part is figuring out the limits... have you figured it out?

OpenStudy (mony01):

is it 0 to 80

terenzreignz (terenzreignz):

0 to somewhere, surely, but where does that tangent line touch the x-axis? Hint: set y = 0, and solve for x.

OpenStudy (mony01):

5x^2=0

terenzreignz (terenzreignz):

No, not the parabola, silly ^_^ the TANGENT LINE :D

OpenStudy (mony01):

x=20

OpenStudy (mony01):

x=2 my bad

terenzreignz (terenzreignz):

Okay, so x runs from 0 to 2, now just integrate the parabola with those limits.

OpenStudy (mony01):

what do you mean?

terenzreignz (terenzreignz):

Integrate 5x^2, with limits from 0 to 2

OpenStudy (mony01):

is it 40/3

terenzreignz (terenzreignz):

Okay, so that's the area of the black shaded part, what about the rest of the yellow area?

OpenStudy (mony01):

do we have to get other limits?

terenzreignz (terenzreignz):

Not only other limits, but another function, too. You notice that this is bounded by the parabola and the LINE, and not the x-axis, so...

OpenStudy (mony01):

true so what other function do we need to use?

terenzreignz (terenzreignz):

Well we can certainly take the area under the parabola, but we must take AWAY the area under the line, any ideas?

OpenStudy (mony01):

ummmm not really?

terenzreignz (terenzreignz):

Take the area under the parabola, which is its integral, and then from that, subtract the area under the line. So what you get is the parabola minus the line: 5x^2 - 40x + 80 Now integrate this... with limits... .... what are the limits? It should start from x = 2 to...?

OpenStudy (mony01):

2 to 80?

terenzreignz (terenzreignz):

No.... you want to get where the line intersects the parabola, and it intersect the parabola at the point where x = ?

OpenStudy (mony01):

5?? sorry im not sure

OpenStudy (mony01):

is it 4?

terenzreignz (terenzreignz):

It's 4, yes. So integrate that new function from 2 to 4

OpenStudy (mony01):

is it 40/3

terenzreignz (terenzreignz):

Yup. So just add that to the first integral, and that's your answer.

OpenStudy (mony01):

so its 80/3

OpenStudy (anonymous):

\[ \int_0^{80} \left(\left(\frac{y}{40}+2\right)-\frac{\sqrt{y}}{\sqrt{5}}\right) \, dy=\frac{1}{40} \left[-\frac{1}{3} 16 \sqrt {5} y^{3/2}+\frac{y^2}{2}+80 y\right]_0^{80}=\frac{80}3 \]

OpenStudy (anonymous):

Here is the way you did it with @terenzreignz \[ \int_0^2 5 x^2 \, dx+ \int_2^4 \left(5 x^2-40 x+80\right) \, dx=\frac{80}{3} \]